# MAVIL: Design of a Multidimensional Assessment of Visual Data Literacy and its Application in a Representative Survey

### v2

@@ Abstract
The ability to read, interpret, and critique data visualizations has mainly been assessed using data visualization tasks like value retrieval. Although evidence on different facets of Visual Data Literacy (VDL) is well established in visualization research and includes numeracy, graph familiarity, or aesthetic elements, they have not been sufficiently considered in ability assessments. Here, VDL is approached as a multidimensional ability whose facets can be partially assessed with self-perceived ability ratings. We introduce an assessment in which VDL is deconstructed as a process of understanding, in reference to frameworks from the learning sciences. MAVIL (Multidimensional Assessment of Visual Data Literacy) comprises six ability dimensions: Comprehending, Decoding, Aestheticizing, Critiquing, Reading, and Contextualizing. MAVIL was designed for general audiences, with validation supported by a survey (N=438) representative of a German-speaking country’s age groups (18-74 years) and male-female gender split. The survey mirrors the population’s VDL—encompassing self-assessment, performance measures and open-ended visualization critique—and shows the understanding of two climate data visualizations, a line and bar chart. For instance, we found that about a quarter have self-assessed deficits in comprehending simple data units, roughly one in five people feel unfamiliar with each displayed chart type and that self-assessed numeracy relates significantly to data reading performance (p=$0.0004$).
@@ 1 Motivation
Visual Data Literacy (VDL) is the human ability to read, interpret, and critique visual data representations [78]. Even though insights into different dimensions of this ability are well established in visualization research and include numeracy [69], graph familiarity [31], or aesthetic elements [70], they remain underutilized in ability assessments. We introduce MAVIL (Multidimensional Assessment of VDL), an approach that integrates self-perceived ability ratings and provides different and more nuanced insights into the understanding of tested visualizations. Through qualitative pilots (N=$8$) and survey validation (N=$438$), we demonstrate its potential for providing insights into VDL facets.
A common assessment strategy for VDL relies on data visualization tasks like value retrieval. Exemplary of such assessment structures are the Visualization Literacy Assessment Test (VLAT) [78], its briefer version, the Mini-VLAT [67], or the Critical Thinking Assessment for Literacy in Visualizations (CALVI) [34]. MAVIL complements existing approaches by incorporating self-perceived abilities, enabling an assessment of VDL that considers both task-based performance and how visualization readers perceive their own skills. By incorporating self-assessment ratings, MAVIL captures nuances that task-based tests may miss, such as how trustworthy a visualization seems or self-indicated difficulties linked to specific graph elements. MAVIL is composed of six ability dimensions:
@@
  • Critiquing: Visualization Criticism,
@@
Designed for general audiences, MAVIL addresses some limitations of prior assessments that strongly focused on expert audiences [31, 40] or specific groups, such as students [13, 30]. Given the increasing need for data literacy—and not everyone being equally data-savvy [52]—our assessment is designed to engage a broad audience, including people of varying ages and without expertise in data visualization.
Here, we showcase the design of MAVIL and the analysis of its implementation in a survey with representative scope for a German-speaking country regarding age groups (18-74 years) and the male-female gender split within them. The survey evaluation illustrates how MAVIL can reveal specific ability dimensions, such as perceived deficits, of visualization readers in a given context. Moreover, it highlights how respondents perceived two tested visualizations, a line chart and bar chart on climate data, chosen for their simplicity, familiarity, and importance in conveying critical information [60, 66, 73].
We found that $48\%$ of respondents (ages 18-74) in the surveyed country make mistakes with the simple line and bar chart. About a quarter of respondents found simple data units, e.g., millimeters, in the surveyed charts incomprehensible. Further, about $12\%$ felt unfamiliar with both displayed graph types. Concerning aesthetic perception, there was a tendency to find the visualization pretty or well-designed and simultaneously rate it as trustworthy. The assessment also included optional text fields for visualization criticism, in which $37\%$ of survey respondents reported at least one point of critique. The qualitative evaluation of text inputs revealed not only respondents’ critical abilities but also which visualization elements can irritate. In our survey, respondents’ self-assessed numeracy level relates significantly (p=$0.004$) to data reading performance and all other dimensions of VDL proposed in this paper.
Ability Dimension
in MAVIL
@@
Summarize visualized content,
recall information [7, 40]
@@
$76\%$ feel confident in
summarizing
@@
Examine color usage, analyze
graph elements [70, 31]
@@
Decide upon trustworthiness,
judge visual appeal [18, 41]
@@
Perceived beauty was linked
to trust
@@
Articulate criticism, make
design suggestions [10, 17]
@@
Solve data reading tasks
[78, 67]
@@
$48\%$ made errors in data
reading tasks
@@
Utilize statistical skill, put
visualization in context [81, 9]
@@
Self-assessed numeracy
correlated with data reading
Table 1: Visual Data Literacy dimensions in MAVIL, developed from either a viewer-focused (top-down) or visualization-focused (bottom-up) perspective. Each dimension is illustrated by example activities and aligned with a corresponding Facet of Understanding [80]. Additionally, we share insights from our representative survey where we tested MAVIL, which is evaluated in detail in Section 5.
@@ 2 Background
Visual Data Literacy (VDL) has been described as the ability to read, interpret, and critically deal with data visualizations [78]. We developed MAVIL’s ability dimensions through a theory-driven approach, drawing on relevant studies and definitions, and combining a bipartite framework with learning sciences concepts that define understanding as a multifaceted process. Creating an assessment that accounts for the multi-dimensional composition of VDL was deemed necessary because current assessments are dominantly task-based, and therefore, they are perspectively limited in their ability display. We highlight studies using self-perceived abilities in relation to performance, arguing that such ratings can offer unique insights into VDL assessments beyond established methods.
2.1 Visual Data Literacy as a Multidimensional Ability
Visual Literacy has been explored as an intellectual and social phenomenon since the 1960s, with early contributions emerging in fields such as education, communication, visual arts, and psychology [58], addressing both audience-related factors and properties of visual design. In MAVIL, we adopt a bipartite framework, that covers top-down (representing the visualization reader) and bottom-up (representing the visualization) aspects of Visual Literacy. Additionally, we draw on insights from the learning sciences, which break complex processes, such as understanding, down into sub-skills.
2.1.1 Considering Viewer and Visualization Perspectives
@@
Visual Literacy can be viewed through a bipartite perspective, distinguishing between top-down and bottom-up aspects [42]. Top-down characteristics take the perspective of visualization readers and are shaped by factors such as prior beliefs, demographic characteristics, and personal skills. Here, numeracy, for instance, has been identified as a critical skill influencing data reading accuracy and knowledge gains [81]. In contrast, bottom-up characteristics focus on the visualization itself, including its content or aesthetic structure [70]. We argue that when assessing Visual Data Literacy, it is valuable to include dimensions that capture the influence of visualization design on audience responses. Visualization techniques, such as color choices or the depiction of uncertainty, have been shown to directly influence how audiences retrieve information and take decisions [61, 79, 25].
Therefore, MAVIL incorporates both top-down and bottom-up aspects by aligning ability dimensions with either a viewer or visualization focus, as shown in Table 1. For example, Contextualizing (Dimension 6) represents top-down characteristics, and Decoding (Dimension 2) relates to bottom-up characteristics. For a full description of dimensions refer to Table 1. We further expanded this two-level approach by integrating frameworks that treat understanding as a multifaceted construct.
2.1.2 Deconstructing Visual Data Literacy into Ability Facets
@@
Insights from learning sciences, such as the framework The Six Facets of Understanding [80], break the process of understanding down into different ability categories: explaining, interpreting, applying, empathizing, having perspective, and self-knowledge. Accordingly, a respondent who can summarize the content of a data visualization (MAVIL Q1c, for an overview of questions, see Figure 1) would be able to explain. If a respondent solves a data reading task (Q5a-Q5d), this is an act of interpretation. Deciding upon the trustworthiness of a visualization (Q3d) requires empathizing with its creator’s intent. Someone who can contextualize displayed data (Q6b) can apply. Critiquing or suggesting design changes (Q4b/Q4c) demonstrates perspective-taking. Self-assessed numeracy (Q6b) or interest in climate change (Q6a) means applying existing skills or knowledge.
@@
Similarly, Bloom’s Taxonomy of Educational Objectives for Knowledge-Based Goals [4] deconstructs knowledge and learning into ability levels across cognitive, affective, and psychomotor domains, which are most commonly used for cognitive intents [53]. According to this framework, memorizing is followed by understanding, application, then analysis, evaluation, and creation. Applied to Visual Data Literacy, a respondent who recalls the information of a visualization (Q1c) remembers it. Someone who integrates information into their prior knowledge (Q1b) can understand. Solving a data reading task (Q5a-Q5d) demonstrates application. Examining color encoding (Q2b) involves analysis. Justifying a point of critique (Q4d/Q4e), represents evaluation. A respondent who suggests a design change (Q4b/Q4c) would be able to create and develop a visualization further. Bloom’s Taxonomy has, for example, been adapted to assess literacy with parallel coordinates plots [68].
Following these approaches from the learning sciences, we built MAVIL dimensions by linking facets of understanding to well-established Visual Data Literacy ability examples, as exemplified in Table 1.
2.2 Limitations of Prior Literacy Assessments
@@
Prior literacy assessments of data visualizations often rely on task-based approaches, focus on a single data visualization type, or target specific audiences. Examples include treemap and parallel coordinate plot literacy tests [30, 27, 28]. Several studies addressed educational contexts, such as qualitative principles for assessing visual literacy in schools [13], while others examine expert understanding of climate data visualizations [31]. Despite these efforts, the utilized measurement approaches remain largely one-dimensional. Moreover, as noted in prior critiques, general audiences with varying ability levels are still underrepresented in visualization research [71, 6, 5].
@@
In 2017, Lee et al. [78] introduced a Visualization Literacy Assessment Test (VLAT), a $53$-item multiple-choice test covering $12$ data visualizations. It is specialized in measuring performance for simple visualization tasks like value retrieval and has been implemented or adapted multiple times [62, 43]. While effective for measuring task-based performance, VLAT’s reliance on correct/incorrect responses limits its ability to capture broader dimensions of Visual Data Literacy, including insights into where deficits may lay and what visualization readers perceive as challenging. VLAT was recently condensed into the Mini-VLAT [67] to improve accessibility by reducing its original 2020+ minute duration. While this may improve accessibility, it still emphasizes numerical accuracy over other skills. The Critical Thinking Assessment for Literacy in Visualizations (CALVI) [34] extends VLAT by shifting its assessment focus on critical ability as part of Visualization Literacy. Comprising 45 items across nine chart types with intentional design flaws, CALVI meaningfully expands the concept of ability but remains focused on task performance. Other assessments ask respondents to construct visualizations themselves, such as in the AVEC assessment [35] or studies on parallel coordinate plot literacy [68], yet all of these remain mainly performance-based.
@@ 2.3 Self-Perception in Ability Assessments
Recently, PREVis, a scale for perceived readability in data visualizations, was introduced by Cabouat et al.[12] and exemplifies how self-assessment may provide insights into literacy. Using Likert scales, it measures four facets of perceived readability: understanding, visualization layout, data value readability, and perception of data patterns. PREVis is built on a strong theoretical foundation and provides a valuable entry point for respondents to self-assess the aspect of perceived readability [12], which we consider part of their visualization understanding. MAVIL overlaps with some PREVis aspects, such as perceived understanding (see MAVIL Q1a in Figure 1) and identifying visual elements (Q2a), but provides additional insights. For instance, our survey, detailed in Section 5, found that $22\%$ of respondents struggled to comprehend "millimeters", and $28\%$ found "millions of tons" unclear (Q2a). PREVis would not capture this hurdle to comprehension.
@@
Self-assessment assumes that human traits and skills are quantifiable and measurable [77] and can complement task-based tests to support more nuanced approaches to measuring abilities [1, 32]. In our study, self-assessment captures how visualization readers perceive facets of their Visual Data Literacy. We hypothesize that these self-perception ratings can reveal nuanced insights into the ability and may correlate with performance in data reading tasks. Similar approaches have been widely applied in psychology, education, and health research: self-concepts of talent have been compared to performance [3], self-perceptions have been linked to academic achievement [76, 36], perceived coping abilities have been shown to predict smoking cessation outcomes [63]. Further, self-perception is recognized as a survey approach in political science, i.e., to gain insights into respondents’ stances and beliefs [59, 50].
@@
While self-perceived ability ratings represent a different measurement type than task-based performance metrics and are not as "objective" or precisely quantifiable, they can still be predictive [2]. Rather than replacing established assessment tools like VLAT [78], we aim to integrate self-perceived ability ratings to enhance Visual Data Literacy assessments. Beyond measuring what people can do, self-assessment sheds light on how respondents perceive their interactions with visualizations. In literacy assessments, factors related to the latter, such as how personal beliefs, trust, or emotional responses influence information retrieval [21, 33], remain underexplored.
3 MAVIL: A Multidimensional VDL Assessment
MAVIL takes a multidimensional approach to Visual Data Literacy assessment, ensuring that all developed questions, including ones on self-perceived ability, are grounded in theoretical frameworks and empirical validation. Four out of six ability dimensions (Dim1-Dim3, Dim6) rely on self-perceived ability ratings, such as Comprehending with Q1a: "I think I understood the visualization". These ratings use five-point Likert scales, indicating levels of Visual Literacy or likelihood of correct interpretation. The according questions are based on a selection of literacy theories and current studies on data visualization understanding. One dimension (Dim5) covers low-level data visualization tasks, which relate to the tested visualization examples, and there are also optional text inputs (Dim4) for visualization critique. An overview of MAVIL’s ability dimensions and question formats is provided in Figure 1, with the dimensions outlined below in the order they were assessed, without implying a hierarchy.

### v3

# A Multidimensional Assessment Method for Situated Visualization Understanding (MdamV)

@@ Abstract
How audiences read, interpret, and critique data visualizations is mainly assessed through performance tests featuring tasks like value retrieval. Yet, other factors shown to shape visualization understanding, such as numeracy, graph familiarity, and aesthetic perception, remain underrepresented in existing instruments. To address this, we design and test a Multidimensional Assessment Method of Situated Visualization Understanding (MdamV). This method integrates task-based measures with self-perceived ability ratings and open-ended critique, applied directly to the visualizations being read. Grounded in learning sciences frameworks that view understanding as a multifaceted process, MdamV spans six dimensions: Comprehending, Decoding, Aestheticizing, Critiquing, Reading, and Contextualizing. Validation was supported by a survey (N=438) representative of Austria’s population (ages 18-74, male/female split), using a line chart and a bar chart on climate data. Findings show, for example, that about a quarter of respondents indicate deficits in comprehending simple data units, roughly one in five people felt unfamiliar with each chart type, and self-assessed numeracy was significantly related to data reading performance (p=$0.0004$). Overall, the evaluation of MdamV demonstrates the value of assessing visualization understanding beyond performance, framing it as a situated process tied to particular visualizations.
@@ 1 Motivation
When interpreting data visualizations, audiences draw on a range of related abilities to read, interpret, and critically engage with them [85, 11]. This range can be referred to as visual data literacy and has typically been assessed through performance tests. Well-known examples include the Visualization Literacy Assessment Test (VLAT) [85], its briefer version, the Mini-VLAT [73], or the Critical Thinking Assessment for Literacy in Visualizations (CALVI) [36], all of which focus on evaluating correct responses to tasks, such as value retrieval. While these instruments are useful for treating literacy as a measurable ability, they fail to capture visualization understanding beyond a strong focus on quantifying test takers’ accuracy in selected-response items. Research has established that facets like graph familiarity [33] and topic knowledge [10] can shape reading accuracy, and shows that literacy is often approached as a multi-faceted set of abilities [55, 56, 74]. Still, such insights remain underutilized in ability assessments.
To address this gap, we frame visualization understanding as the application of visual data literacy in context-specific settings, meaning, in direct relation to the visualizations used in the study. We present a Multidimensional Assessment Method of Situated Visualization Understanding (MdamV) that was applied to static climate data visualizations. MdamV was developed in alignment with learning sciences theory on understanding as a multifaceted process. It was then validated through qualitative pilots with eight participants and a survey (N=$438$) representative of the Austrian population’s age groups from 18 to 74 years and their male/female gender split. Our method combines task-based measures with self-perceived ability ratings and open-ended critique. This combination captures not only whether respondents can complete tasks correctly, but also how they perceive their own understanding and indicate difficulties with specific visualization features. Therefore, it reveals nuances that established assessments may miss, such as perceived trustworthiness or challenges with decoding color. MdamV is composed of six assessment dimensions:
@@
  • Critiquing: Visualization Critique,
@@
Designed for general audiences, MdamV addresses some limitations of prior assessments that strongly focused on expert audiences [33, 43] or specific groups, such as students [14, 32]. Given the increasing need for data literacy—and not everyone being equally data-savvy [58]—our survey targeted a broad audience, including people of varying ages and without expertise in data visualization.
Here, we showcase the design of MdamV and report on its application in a survey representative of Austria’s population (ages 18-74, male/female split). Our results demonstrate that a multidimensional approach can yield informative insights into audience understanding beyond visualization task performance. They also show how respondents engaged with two tested visualizations, a line chart and a bar chart on climate data, chosen for their simplicity, familiarity, and importance in conveying critical information [66, 72, 80]. As MdamV is evaluated with two specific visualizations, it should not be seen as a ready-to-use instrument but rather a method with the potential to be adapted and applied in different contexts.
In our representative survey, we found that $48\%$ of respondents in Austria make mistakes with the simple line and bar chart. About a quarter of respondents found simple data units, e.g., millimeters, in the surveyed charts incomprehensible. Further, about $12\%$ felt unfamiliar with both displayed chart types. Concerning aesthetic perception, there was a tendency to find the visualization pretty or well-designed and simultaneously rate it as trustworthy. In optional text fields, $37\%$ of survey respondents reported at least one point of critique, offering qualitative insights into both their critical engagement and the visualization elements they found irritating. In our survey, respondents’ self-assessed numeracy level relates significantly (p=$0.004$) to data reading performance.
Assessment Dimension
in MdamV
@@
Summarize visualized content, recall
information [8, 43]
@@
$76\%$ feel confident in summa-
rizing
@@
Examine color usage, analyze graph
elements [77, 33]
@@
Decide upon trustworthiness, judge
visual appeal [19, 44]
@@
Perceived beauty was linked to
trust
@@
Articulate critical thought, recognize
the meaning of construction[11, 18]
@@
Solve data reading tasks [85, 73]
@@
$48\%$ made errors in reading
tasks
@@
Utilize statistical skill, put visual-
ization in context [88, 10]
@@
Self-assessed numeracy corre-
lated with data reading
Table 1: Assessment dimensions covered in MdamV, developed from either a viewer-focused (top-down) or visualization-focused (bottom-up) perspective. Each dimension is illustrated by example activities and aligned with a Facet of Understanding [87]. Additionally, we share insights from our representative survey where we applied MdamV to two climate data visualizations. The results are analyzed in detail in Section 5.
@@ 2 Background
Through a theory-driven approach, we developed six different assessment dimensions for visualization understanding, drawing on prior work related to visual data literacy. We frame visualization understanding as the situated application of visual data literacy: a process in which a range of abilities come together to understand a particular visualization. Therefore, we consider both the reader’s perspective and the characteristics of the visualization [45, 77, 64]. Building on this, we align our assessment dimensions with learning sciences frameworks that define understanding as "a family of related abilities" [87]. Creating a multidimensional approach was deemed necessary because current assessments are dominantly task-based, and therefore, they are perspectively limited in their ability display. We highlight studies using self-perceived abilities in relation to performance, arguing that such ratings can offer unique insights into visualization understanding beyond established tests.
2.1 Visual Data Literacy: Viewer and Visualization Perspective
@@
Visual literacy has been explored as an intellectual and social phenomenon since the 1960s, with early contributions emerging in fields such as education, communication, visual arts, and psychology [64], addressing both audience-related factors and properties of visual design. Accordingly, it can be viewed through a bipartite perspective, distinguishing between top-down and bottom-up aspects [45]. Top-down characteristics take the perspective of visualization readers and are shaped by factors such as prior beliefs, demographic characteristics, and personal skills. Here, numeracy, for instance, has been identified as a critical skill influencing data reading accuracy and knowledge gains [88]. In contrast, bottom-up characteristics focus on the visualization itself, including its content or aesthetic structure [77]. We argue that it is valuable to capture visualization understanding through dimensions that consider the influence of visualization design on audience responses. Visualization techniques, such as color choices or the depiction of uncertainty, have been shown to directly influence how audiences retrieve information and make decisions [67, 86, 26].
Since we recognize both top-down and bottom-up aspects, we align our assessment dimensions with either a viewer- or visualization-centered focus, as shown in Table 1. For example, questions on Contextualizing represent a top-down perspective and include viewers’ numeracy and topic interest, while questions on Decoding reflect a bottom-up perspective by referring to color encoding and chart type.
2.2 Deconstructing Visualization Understanding into Ability Facets
@@
Insights from learning sciences, such as Wiggins and McTighe’s framework of the Six Facets of Understanding [87], conceptualize understanding as a multifaceted process composed of different abilities: explaining, interpreting, applying, empathizing, having perspective, and self-knowledge. These facets can be aligned with representing different abilities that are applied during visualization understanding. For example, a respondent who can summarize the content of a data visualization (see Q1c in Fig. 1) would be able to explain. If a respondent solves a data reading task (Q5a-d) or translates color encoding (Q2b), this is an act of interpretation. Deciding upon the trustworthiness of a visualization (Q3d) requires empathizing with its creator’s intent. Critiquing or suggesting design changes (Q4b/c) indicates perspective-taking. Someone who can contextualize displayed data (Q6b) can apply, e.g., by drawing on their knowledge of the visualization’s topic. Finally, self-knowledge involves consciously questioning one’s own abilities and limits in understanding. While this could be valuable to assess, it lay beyond the scope of our survey design, e.g., by considering readers’ intrinsic habits of mind and how these shaped their engagement.
Building on these categories from the learning sciences, we aligned assessment dimensions with facets of understanding to illustrate our procedure and to better exemplify visualization understanding as a process involving a multidimensional set of abilities (see Tab. 1).
@@
Another learning sciences framework worth considering is Bloom’s Taxonomy of Educational Objectives for Knowledge-Based Goals [4], which deconstructs learning into ability levels (memorizing, understanding, application, analysis, evaluation, and creation), which are most commonly used for cognitive domains [59]. While our alignment focused on the Six Facets of Understanding, this taxonomy offers a complementary view, for example, by linking the recall of information (Q1c) to memorizing, the salvation of a data reading task (Q5a-d) to application, or the examination of color encoding (Q2b) to analysis. It has already been adapted to assess literacy in contexts such as parallel coordinates plots by Peng et al. [75].
2.3 Limitations of Prior Assessments
@@
Prior assessments related to visual data literacy often rely on task-based approaches, focus on a single data visualization type, or target specific audiences. Examples include treemap and parallel coordinate plot literacy tests [32, 29, 30]. Several studies addressed educational contexts, such as qualitative principles for assessing visual literacy in schools [14], while others examine expert understanding of climate data visualizations [33]. Despite these efforts, the utilized measurement approaches remain largely one-dimensional. Moreover, as noted in prior critiques, general audiences with varying ability levels are still underrepresented in this line of research [78, 7, 6].
@@
In 2017, Lee et al. [85] introduced a Visualization Literacy Assessment Test (VLAT), a $53$-item multiple-choice test covering $12$ data visualizations. It specializes in measuring performance for simple visualization tasks like value retrieval and has been implemented or adapted multiple times [68, 46]. While effective for measuring task-based performance, VLAT’s reliance on correct/incorrect responses limits its ability to capture other facets of visualization understanding, such as where deficits may lay and what visualization readers perceive as challenging. VLAT was recently condensed into the Mini-VLAT [73] to improve accessibility by reducing its original 2020+ minute duration. While this may improve accessibility, it still emphasizes numerical accuracy over other abilities. The Critical Thinking Assessment for Literacy in Visualizations (CALVI) [36] builds on this approach by incorporating critical ability as part of visualization literacy. Comprising 45 items across nine chart types with intentional design flaws, CALVI meaningfully expands the assessment approach but remains focused on task performance. In other approaches, respondents are asked to construct visualizations, such as in the AVEC assessment [37] or studies on parallel coordinate plot literacy [75], yet they too are primarily based on task performance.
@@ 2.4 Self-Perception in Ability Assessments
Recently, PREVis, a scale for perceived readability in data visualizations, was introduced by Cabouat et al. [13] and exemplifies how self-assessment may provide insights into visualization understanding. Using Likert scales, it measures four facets of perceived readability, including visualization layout, and data value readability. PREVis provides a valuable entry point for respondents to self-assess the aspect of perceived readability [13]. MdamV overlaps with some PREVis aspects, such as perceived understanding (see Q1a in Fig. 1) and identifying visual elements (Q2a), but provides additional insights. For instance, our survey, detailed in Section 5, found that $22\%$ of respondents indicated comprehension issues with millimeters, and $28\%$ with millions of tons (Q2a). This level of detail, tied to tested visualizations and their features, highlights comprehension challenges that PREVis would not capture.
@@
Self-assessment assumes that human traits and skills are quantifiable and measurable [84] and can complement task-based tests to support more nuanced approaches to measuring abilities [1, 34]. In our method, self-assessment measures capture how visualization readers perceive facets of their visualization understanding. We hypothesize that these ratings can reveal nuanced insights into the respondents’ set of abilities and may correlate with performance in data reading tasks. Similar approaches have been widely applied in psychology, education, and health research: self-concepts of talent have been compared to performance [3], self-perceptions have been linked to academic achievement [83, 38], perceived coping abilities have been shown to predict smoking cessation outcomes [69]. Further, self-perception is recognized as a survey approach in political science, i.e., to gain insights into respondents’ stances and beliefs [65, 54].
@@
While self-perceived ability ratings differ from performance metrics and are not as "objective" or precisely quantifiable, they can still be predictive [2]. Rather than replacing task-based instruments like VLAT [85], we integrate self-perceived ability ratings to broaden how visual data literacy is being approached. Beyond measuring what people can do, self-assessment sheds light on how respondents perceive their interactions with visualizations. In literacy assessments, factors related to the latter, such as how personal beliefs, trust, or emotional responses influence information retrieval, remain underexplored [22, 35].
3 MdamV: A Multidimensional Assessment Method for Situated Visualization Understanding
With MdamV (Multidimensional Assessment Method), we suggest six assessment dimensions (see Fig. 1) that we constructed for visualization understanding and evaluated in a survey. Four out of six dimensions (Comprehending, Decoding, Aestheticizing, Contextualizing) rely on self-perceived ability ratings (see Sec. 2.4). One dimension (Reading) covers low-level visualization tasks, while another offers optional text inputs (Critiquing) for open-ended critique. The dimensions are outlined below in the order they were assessed, without implying a hierarchy. Similar to how Wiggins and McTighe [87] describe for their framework on understanding (see Sec. 2.2), we do not claim our proposed dimensions to be the only way of capturing understanding. Instead, we aim to break visualization understanding into different facets, making underexplored abilities in prior assessments (see Sec. 2.3) more visible.